Solve the following triangles. Identify the case # in each given
triangle.

5.) Given parts:
A= 78°48' b= 726 c= 938
1
Expert's answer
2020-11-15T18:16:58-0500

"a = \sqrt{b^2 +c^2 -2b*c*cos\angle{A}}"

"a = \sqrt{726^2 +938^2 -2*726*938*cos78\degree 48'} = 1650.168"

"a / sin\angle A = b/sin\angle B"

"sin\angle B = b*sin\angle A / a = 726*sin78.8\degree / 1650.168 = 0.432"

if "\angle B > 90\degree," when "\angle C < 11\degree 12'" and side the smallest, but the smallest side is b,

therefore "\angle B < 90\degree,"

"\angle B = 28.59\degree = 28\degree36'"

"\angle C = 180\degree -\angle A -\angle C = 180\degree - 28\degree36' - 78\degree 48'"

"\angle C = 72\degree36'"

answer :

"a =1650.168\\\angle C = 72\degree36'\\\angle B = 28\degree36'"